Any help would be very kind. "Write the equation of the line. Slope=0, passes through (-4, -5)" thank you.
Answer by sofiyac(983)

the equation for this type of problem where you are given the slope and a point is
y-y1=m(x-x1) where m is the slope and (x1,y1) is the given point. Just plug everything in.
y-(-5)=0(x-(-4)) anything times 0 is 0 so right side automatically becomes 0
y+5=0 subtract 5 from each side
y=-5 that's your equations, it's basically just a straight horizontal line through y=-5